OpenStudy (anonymous):

whats the equation? 3over5; y intercept (0,-5)

OpenStudy (anonymous):

I'm assuming you are looking for y-intercept form (y= mx +b)? 3/5 is your slope. So you will want to find b for the equation. -5 = 3/5(0) + b. -5 = 0 + b. b = -5. So the equation would be \[y = \frac{ 3 }{ 5 }x - 5\]

OpenStudy (anonymous):

brandi rocks I had no idea what that question was doin

OpenStudy (anonymous):

That is amazing thank you...